# On the distance between mean and geometric median in high dimensions

### v2

@@ A.1 Proofs
$\operatorname{E}\left[\,A_{p,i}^{2}\,\right]=\operatorname{E}\left[\,((X_{i}-m_{p,i})^{2})^{2-q/2}|X_{i}-m_{p,i}|^{q}1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}\,\right]\leq\\ p^{2-q/2}\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq p^{2-q/2}\tilde{C}$.
Thus, in any case, $\operatorname{Var}\left[\,A_{p,i}\,\right]\leq\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\tilde{C}\cdot\max(1,p^{2-q/2})$.
Let $i\in\mathbb{N}$. By $M$-dependence, $\operatorname{Cov}(A_{p,i},A_{p,j})=0$ for all but $2M+1$ indices $j$. This fact together with $|\operatorname{Cov}(X,Y)|\leq\sqrt{\operatorname{Var}\left[\,X\,\right]\operatorname{Var}\left[\,Y\,\right]}$ gives
@@
Combining (30) with (29) in (28), we have shown
@@
in probability as $p\to\infty$. Further, note
@@
A.2 Further analysis of the simulation results

### v3

@@ A.5 Additional Lemmas
$\operatorname{E}\left[\,A_{p,i}^{2}\,\right]=\operatorname{E}\left[\,((X_{i}-m_{p,i})^{2})^{2-q/2}|X_{i}-m_{p,i}|^{q}1_{\{(X_{i}-m_{p,i})^{2}\leq p\}}\,\right]\\ \leq p^{2-q/2}\operatorname{E}\left[\,|X_{i}-m_{p,i}|^{q}\,\right]\leq p^{2-q/2}\tilde{C}$.
Thus, in any case $\operatorname{Var}\left[\,A_{p,i}\,\right]\leq\operatorname{E}\left[\,A_{p,i}^{2}\,\right]\leq\tilde{C}\cdot\max(1,p^{2-q/2})$.
Let $i\in\mathbb{N}$. By $M$-dependence, $\operatorname{Cov}(A_{p,i},A_{p,j})=0$ for all but $2M+1$ indices $j$. This fact together with $|\operatorname{Cov}(X,Y)|\leq\sqrt{\operatorname{Var}\left[\,X\,\right]\operatorname{Var}\left[\,Y\,\right]}$ gives:
@@
Combining (30) with (29) in (28), we have shown that
@@
in probability as $p\to\infty$. Further, note that
@@
Appendix B Further analysis of the simulation results